CielBleu:

Determine the magnitude of the difference of the following vectors: u - v u = 7i - 12j v = -2i + j

TheSmartOne:

u = <u1,u2> v = <v1,v2> u-v = <u1-v1, u2-v2> Magnitude = \(\sqrt{ (u_1-v_1)^2 + (u_2-v_2)^2}\)